The radius of the light circle observed by a fish at a depth of 12 meter is (refractive index of water = 4/3)
(A) 736 (B) 7
36
(C) 536 (D) 54

Grade:

2 Answers

Arun
3 years ago
sinic=μ1=r2+h2r
Using h=12 cm, μ=4/3
We get r=736 cm.
 
Hope it helps
In case of any difficulty, please feel free to contact again
 
Thanks and regards
Arun
Vikas TU
3 years ago
sinθ = 1/μ = 3/4 
or tanθc=3/ √(16−9) = 3/√7
= R/15 
impliesR=(36)/(sqrt7)`cm................................